form the differential equation for this
y=ae^x+bln|y|+c+d
1
Expert's answer
2012-10-30T12:09:26-0400
Derivation by dy and dx gives us
d(y-blny)/dy*dy=d(ae^x+c+d)/dx*dx
(1-1/y)dy=ae^xdx
y*y'-y'=a*e^x*y will be an equation.